There are 630 dishes that need to be rinsed. John can rinse them in 70 minutes by himself. It will take his friend Bob 105 minutes to rinse these dishes. How long will it take them if they rinse these 630 dishes together?
Answer by nyc_function(2741)

70 minutes = 7/6 hours.

105 minutes = 7/4 hours.

[(7/6)(7/4)]/[(7/6) + 7/4)] = time working together.

Simplify the complex fraction to find your answer.